OpenStudy (anonymous):

Help with 3 math questions? :3 Medal! ^.^

OpenStudy (anonymous):

The graph below shows this system of linear equations.

OpenStudy (anonymous):

What is the solution to the system of linear equations? (4, 4) (4, 5) (0, -3) (0, 1)

OpenStudy (anonymous):

The solution is the point of intersection on the graph. Do you see it?

OpenStudy (anonymous):

That's for question 2

OpenStudy (anonymous):

you can use elimination to solve it ?

OpenStudy (anonymous):

The equations fit the picture you've given. You can use elimination or substitution. Probably substitution is easiest for this problem since you've got y= something already. To solve: x+1=2x-3 Get the x values on one side and the numbers on the other.

OpenStudy (anonymous):

x+1=2x-3 1+3=2x-x 4=x Since you now know x, substitute 4 into the equations above. You'll find y.

OpenStudy (anonymous):

4+1=y 2(4)-3=y 5=y 8-3=y 5=y solution is (4,5)

OpenStudy (the_fizicx99):

Its unnecessary to solve using the elimination process.. just look at when both lines intersect. (4,5) .. they're already in slope intercept form.